OpenStudy (anonymous):

make n the subject of the formula: A = 5mn + pn

OpenStudy (anonymous):

first divide throughout by n

OpenStudy (anonymous):

so A=5mn+pn A/n=5mn/n+pn/n A/n=5m+p

OpenStudy (anonymous):

then divide throughout by A so A/n /A =(5m+p)/A 1/n=(5m+p)/A

OpenStudy (anonymous):

then 1 divided by 1/n = 1 divided by (5m+p)/A so n=A/(5m+p)

OpenStudy (anonymous):

thank you !!!

OpenStudy (anonymous):

welcome